OpenStudy (anonymous):

1. What is the domain of the function? Y=V3x+3 (V is the square root)

OpenStudy (anonymous):

we cannot take the square root of any negative values so solve the inequality\[3x+3\ge0\]

OpenStudy (anonymous):

The answer is either a 1 or a -1 im just not totally sure

OpenStudy (anonymous):

is the function \[\sqrt{3x+3} or \sqrt{3x} +3\]

OpenStudy (anonymous):

Its the first one

OpenStudy (anonymous):

if the first it would be -1. 3(-1)+3=0, and 3(1)+3>0, 6

OpenStudy (anonymous):

So the domain of the function is \[[-1, \infty)\]

OpenStudy (anonymous):

would the answer be x\[\ge-1\]

OpenStudy (anonymous):

That, or the above. Depends how your teacher prefers your answer to be expressed.

OpenStudy (anonymous):

The answers are all expressed like the way i typed